# RuCCoD: Towards Automated ICD Coding in Russian

### v1

@@ Appendix D Entity Linking Results
RuCCoN (Nesterov et al., 2022a) is a manually annotated corpus of clinical records in Russian. It contains 16,028 mentions linked to 2,409 unique concepts from the Russian subset of UMLS metathesaurus (Bodenreider, 2004).
@@
The results of cross-terminology entity linking transfer presented in Tab. 8 reveal a few insightful findings related to linking ICD codes.
@@ Appendix E LLM with RAG results
All LLM with RAG experiments were conducted with a temperature setting of 0 for all LLMs and a top-k value of 15 for the number of retrieved entities from similarity search. The LLMs used are specified in Appx. G. For the embedding model, we utilized BERGAMOT. To construct the vector database, we used dictionaries extracted from NEREL-BIO, RUCCON, the ICD dictionary, and the ICD dictionary combined with RuCCoD. The results are presented in Tables 9 and 10 for strict evaluation, and in Tables 11 and 12 for relaxed evaluation.
@@ Appendix H Prompts
NER prompt You will be provided with a text containing diagnoses. Extract the diagnoses from this text. Do not alter the spelling of the diagnoses in the text. Respond only in the format of a list: [’diagnosis1’, diagnosis2’, ...] Text: {text}
Diagnosis selection prompt You will be given a reference diagnosis and a list of diagnoses from a database. Your task is to determine which diagnosis from the database best matches the reference diagnosis. Try to select the diagnosis accurately, paying attention to details. Choose the diagnosis with the highest match in terms of words and meaning. You can only choose from the diagnoses in the list. Pay more attention to the diagnoses at the beginning of the list, as they are more likely to be a better match. Its better to choose a shorter diagnosis than one that includes information not present in the reference diagnosis. In your response, write only the diagnosis number and nothing else. Reference diagnosis: {diagnosis} List of diagnoses from a database: {list}

### v2

@@ Appendix D Entity Linking Results
RuCCoN (Nesterov et al., 2022) is a manually annotated corpus of clinical records in Russian. It contains 16,028 mentions linked to 2,409 unique concepts from the Russian subset of UMLS metathesaurus (Bodenreider, 2004).
@@
The results of cross-terminology entity linking transfer presented in Tab. 7 and Tab. 9 reveal a few insightful findings related to linking ICD codes.
@@ Appendix E LLM with RAG results
All LLM with RAG experiments were conducted with a temperature setting of 0 for all LLMs and a top-k value of 15 for the number of retrieved entities from similarity search. The LLMs used are specified in Appx. G. For the embedding model, we utilized BERGAMOT. To construct the vector database, we used dictionaries extracted from NEREL-BIO, RuCCoN, the ICD dictionary, and the ICD dictionary combined with RuCCoD. The results are presented in Tables 10 and 11 for strict evaluation, and in Tables 12 and 13 for relaxed evaluation.
@@ Appendix H Prompts
NER prompt You will be provided with a text containing diagnoses. Extract the diagnoses from this text. Do not alter the spelling of the diagnoses in the text. Respond only in the format of a list: [’diagnosis1’, ’diagnosis2’, …] Text: {text}
Diagnosis selection prompt You will be given a reference diagnosis and a list of diagnoses from a database. Your task is to determine which diagnosis from the database best matches the reference diagnosis. Try to select the diagnosis accurately, paying attention to details. Choose the diagnosis with the highest match in terms of words and meaning. You can only choose from the diagnoses in the list. Pay more attention to the diagnoses at the beginning of the list, as they are more likely to be a better match. It’s better to choose a shorter diagnosis than one that includes information not present in the reference diagnosis. In your response, write only the diagnosis number and nothing else. Reference diagnosis: {diagnosis} List of diagnoses from a database: {list}